### v1

@@ 3 Results
Figure 2: Percent Reduction for the Waymo ADS in Police-Reported and Any-Injury-Reported Crashes Compared to a Human Benchmark from Scanlon et al. (2023) in Phoenix, San Francisco, and All Locations. Note that Los Angeles data is not plotted because of large error bars. See the results in Table 5
4 Discussion
4.1 Interpretations of Results
4.1.1 Any Property Damage or Injury
The available benchmark crashed vehicle rates for any property damage or injury crashes vary greatly depending on the method of estimation (i.e., underreporting adjustment of crash data vs. observations in NDS) and the driving environment (e.g., SF ride-hailing NDS vs. national average NDS). These observed discrepancies could be due to several factors. As noted in Scanlon et al. (2023), although geographic location is controlled for in the benchmarks used in this paper, other driving environment factors are not. For example, the ride-hailing benchmark from San Francisco in Flannagan et al. (2023) will have some amount of differing driving exposure (e.g., time of day, location/routes) than the overall passenger vehicle populations derived from police report data because the drivers studied in Flannagan et al. (2023) were primarily using their vehicles to offer ride-hailing services. Another possible explanation for this discrepancy is that the underreporting adjustment used to estimate any property damage or injury crashes could also vary by driving environment. The available underreporting estimates are few, and are derived from large, nationally representative surveys that may not capture reporting differences in different driving environments. Last, it is still not clear if the minimum severity for reporting is consistent between different benchmark sources and the ADS fleet. The effect of including or excluding low delta-V crashes can also be observed in the ADS crash data. Excluding vehicle-to-vehicle crashes with low delta-V (less than 1 mph) reduced the ADS crashed vehicle rates by 48% in Phoenix and 52% in San Francisco.
The assumed crash underreporting used in Scanlon et al. (2023) to adjust police-reported data to estimate the any property damage or injury outcome level are subject to uncertainties. When surveying people about underreporting, there is subjective judgment on what is considered a crash. The survey studies used as part of the underreporting adjustment in Blincoe et al. (2022) and subsequently used by Scanlon et al. (2023) were open-ended and asked respondents to report whether they were involved in a crash, without defining what a crash constitutes in terms of vehicle damage. On the NDS side, there can be similar difficulties judging property damage status of crashes. Most NDS will identify most crashes in the data reduction phase based on kinematic triggers and manual review of the sensor and video data. Data reductionists need to make subjective judgements on the severity of these crashes without the benefit of retrospective vehicle damage investigations. Although this NDS-based method may discover most crashes, there are certain types of low impact severity crashes that may be more likely to be missed (e.g., while the equipped vehicle is stationary and thus do not trigger kinematic triggers or those that occur outside the camera views of the instrumentation). Further complicating property damage crash reporting, ADS vehicles are equipped with high cost sensors that may be damaged in even minor crashes. If a property damage value approach is used, ADS vehicles may have higher repair costs, and thus have a higher property damage crash rate for a similar set of benchmark crash with similar impact energy characteristics.
In addition to underreporting uncertainty, as discussed in Scanlon et al. (2023), a stated limitation of the state-wide police report database in California (SWITRS) is that there is no requirement that non-injury police-reported (within a jurisdiction) crashes be reported to SWITRS (the state). This potential incomplete reporting of property damage crashes, even those that meet the reporting requirement, could explain why San Francisco has a lower police-reported crashed vehicle rate than Phoenix, even though the expectation is that there are more crashes in more densely populated areas. Although San Francisco had a lower any property damage or injury crashed vehicle rate compared with Phoenix, San Francisco had the highest any-injury-reported crashed vehicle rate. It is unclear whether this result is because San Francisco indeed has a higher ratio of any-injury-reported to any property damage or injury crashes than other locations, or if this is an artifact of this underreporting that affects SWITRS data specifically.
Because of these discrepancies and difficulties, it is difficult to draw conclusions about the relative performance of an ADS like the Waymo Driver and a benchmark crashed vehicle rate for any property damage or injury crashes. Although the results show that the Waymo ADS has, in general, a lower crashed vehicle rate in any property damage or injury crash outcome group, it is not clear what the appropriate benchmark should be. For example, in San Francisco the any property damage or injury benchmark from Scanlon et al. (2023) is between 10.5 and 17.9 IPMM (depending on the underreporting modeling adjustment methodology), while the ride-hailing driver benchmark in San Francisco reported in Flannagan et al. (2023) is 64.9 IPMM. Given the available data, it’s unclear if this discrepancy is because the underreporting adjustment used in Scanlon et al. (2023) is not representative of San Francisco, whether the SWITRS police report database is underreporting property damage only reports at a higher rate than elsewhere, or if the different characteristics of the driving populations (all drivers vs ride-hail drivers using rental vehicles) are affecting the rates. In order to better understand any property damage or injury crashed vehicle rates, there is a need for metrics or data sources that can more accurately define a more objective lower reporting threshold than exists today.
@@ 4.1.1 Any Property Damage or Injury
In this study, the ADS any property damage or injury crashes were reported in two severity levels: all reported contacts and those that excluded crashes with delta-V less than 1 mph. The availability of accurate sensor data enables accurate reconstructions of ADS crashes to estimate their severity (e.g., delta-V or impact speed). Estimating crash severity is often not possible from most police report databases without detailed reconstructions. The Crash Investigation Sampling System (CISS), maintained by NHTSA, is a nationally representative database where some crashes have an accompanying crash reconstruction based on vehicle damage and scene evidence. This sampled, nationally representative database might be difficult to apply to different driving environments, however. As shown in the benchmarks of Scanlon et al. (2023), crashed vehicle rates can vary widely in different locations. Another potential source of severity estimates in a benchmark is estimating severity from NDS data. Advances in computer vision may enable crash reconstructions from limited vehicle sensor and video data (see for example Campolettano et al. (2023)).
4.1.2 Police-Reported and Any-Injury-Reported
@@
This comparison of police-reported crashed vehicle rates is likely a conservative comparison for several reasons. Private citizens often do not report police-reportable events (M. Davis & Co., 2015). Again, in California, only injury crashes are required to be reported to the state databases, which means some jurisdictions in California may have different PDO police-reported crashes reporting practices. Conversely for ADS crashes, this study conservatively used the “Law Enforcement Investigating?” field of the NHTSA SGO reports because the authors have knowledge that Waymo’s operational policies where it is likely police took information that could be used to file a report or indicated they would file a report. Although the police-reported outcome is often associated with a certain level of property damage, factors other than property damage may influence whether an ADS fleet operator would report collisions to the police. Future research should investigate police reporting differences between businesses (like an ADS fleet operator) and the general population. In some cases, Waymo was not provided with a filed report. Because most of the crashes in this dataset occurred in 2022 and 2023, for which state police-reported databases have not been published at the time of writing, the authors could not verify whether the ADS crashes in this study in the police-reported outcome group indeed were recorded. The benchmark from Scanlon et al. (2023) used for comparison in the police-reported outcome group was recorded in the respective crash databases.
@@
any-injury-reported crashes may also serve as a useful comparison metric between ADS and a benchmark. It has been shown that there is less underreporting in injury crashes compared to crash with no injuries Blincoe et al. (2022). Because the magnitudes of underreporting are smaller, the variance and the impact on the comparisons may also be lower. The injury rates, however, can also suffer from similar surveillance biases than the property damage and police report data related to underreporting. The minimum threshold of an injury is often unclear. The timing of presentation of injury can also affect the results. Injury determination as reported in police reports is often reported on scene when the report is filed. A police report may not reflect injuries that are reported days after a crash. ADS reported crashes are often filed days, weeks, or up to a month after the event, allowing for more time to become aware of alleged injuries from those involved.
4.2 Comparison to Insurance Claims-based Analyses
The benchmarks and ADS data presented in this paper are overall crash involvement rates which do not consider crash contribution of those involved. Di Lillo et al. (2023) compared third party liability property damage and bodily injury insurance claims rate between the Waymo ADS operations and a human benchmark. In this study, only liability claims that were resolved or were likely to be resolved with a payment were included as these types of claims are a proxy for responsibility. Di Lillo et al. (2023) found that the Waymo RO service had a lower property damage and bodily injury claims rate compared to a human benchmark over 3.8 million miles of RO driving. Examining liability claims allows an analysis of the safety impact of a population (here the Waymo RO service) towards other road users in the way the ADS contributes to causing crashes. The results of this study, in contrast, examine the overall crashed vehicle rate regardless of the contribution to the crash’s causation. Indeed, many of the crashes in the RO dataset examined in this paper featured the Waymo ADS parked or stopped appropriately at a traffic control device where the ADS’s behavior had little or no contribution to the crash’s cause. Examining ADS performance from both a crash causation and overall crash rate are important and can provide information on the overall safety impact of an ADS. Di Lillo et al. (2023) reported that Waymo RO operations reduced property damage liability claims by 76% and bodily injury liability claims by 100% (zero observed RO bodily injury liability claims). This magnitude of claims reduction is similar to the magnitude in police-reported (51% to 70% reduction) and any-injury-reported (76% and 90% reduction) reductions reported in this study, which did not consider contribution to causing the crash.
4.3 Considerations for Safety Impact Assessment for Serious Injury and Fatal Crashes
Conclusions regarding serious injury alone, or fatalities alone, require more data because they are subsets of the any-injury-reported outcome group which combines all injury and fatality results into one data category. ADS developers use multiple, complementary safety assurance methods to mitigate risk of high severity outcomes across architectural, behavioral, and operational layers (e.g., platform verification and validation, hazard analysis, see Webb et al. 2020, Favaro et al. 2023a). Traditionally, automotive safety systems have used prospective safety benefit analyses that use test track data and simulations as risk assessment tools and to determine potential high severity benefits. These types of prospective studies have also been performed for ADS. For example, one study that used reconstructions of all fatal crashes in Chandler, AZ found that a level 4 ADS was able to prevent 100% of crash when placed in the initiator role and prevent 82% and mitigate 10% when placed in the responder role (Scanlon et al., 2021).
Furthermore, ADS are designed to be able to follow applicable speed limits on the roads they drive on. Reducing speeding is a core pillar of the Vision Zero and Safe Systems approach to reducing serious injuries and fatalities on roadways. The Safe Systems approach has been adopted as a policy initiative by the U.S. Department of Transportation (United States Department of Transportation, 2022). The results of this study show that the ADS evaluated had a statistically significant lower any-injury-reported crashed vehicle rate. Serious injury and fatalities are a subset of this any-injury-reported benchmark, but no statement on these outcome levels can be made at this time based on this retrospective data. In a Safe Systems approach, it is not a requirement to prove efficacy of countermeasures before they are deployed in the field. Due to the statistical requirements, like the ones presented in Scanlon et al. (2023) or Kalra and Paddock (2016), assessing the safety impact of a system on serious injury or fatal crashes is impractical in many situations. The Safe Systems approach instead uses known causal factors that contribute to a large amount of serious injury and fatal crashes (e.g., excessive speed, or excessive energy transfer) and attempts to control these factors to stay below the tolerances to lead to these outcomes.
As ADS RO miles increase, the likelihood of observing a serious injury or fatality also increases. As discussed in the statistical power analysis of Scanlon et al. (2023), hundreds of millions to billions of miles of VMT would be necessary to detect a statistically significant difference in fatal crashed vehicle rates, depending on the magnitude of the difference in crashed vehicle rate of the ADS and benchmark. The national rate of crashed passenger vehicles involved in fatal crashes while traveling on surface streets is 22.3 incidents per billion miles (IPBM) (Scanlon et al., 2023). Consider a fictive driver who experiences 1 fatal crash after driving 10 million miles. At the time of this fictive fatal crash, the driver’s fatal crashed vehicle rate would be 100 IPBM with 95th percentile confidence intervals of [3, 557] IPBM. Although the point estimate of 100 IPBM is higher than the benchmark, due to the low event count relative to the miles, this fictive driver’s crashed vehicle rate is not statistically different from the benchmark. Consider another fictive driver involved in 1 fatal crash in 100 million VMT, which is a rate of 10 IPBM. The 95th percentile confidence intervals of this fictive driver is [0.3, 56] IPBM, which is still overlapping with the benchmark and not significantly different.
4.4 Limitations
@@
This study has several limitations. The results in this study examine crashed vehicle rates aggregated by operating location and do not attempt to evaluate crashed vehicle rates within subgroups, for example, by crash partner or crash type. As discussed above, the method used for estimating an overall crashed vehicle rate, and particularly underreporting assumptions, affect the ability to draw conclusions about any property damage or injury crashed vehicle rate performance. Past studies that reported ADS crash data have found that ADS vehicles may have different proportions of different crash rates than the general population (Victor et al., 2023).

### v2

@@ RESULTS
Figure 3: Percent Reduction for the Waymo ADS in Police-Reported and Any-Injury-Reported Crashes Compared to a Human Benchmark from Scanlon et al. (2023) in Phoenix, San Francisco, and All Locations. Note that Los Angeles data is not plotted because of large error bars. See the results in the Appendix
DISCUSSION
Comparison to Previous Studies
This study is one of the first to compare RO ADS crash rates to comparable benchmarks in the geographic locations in which ADS currently operate. In addition to comparing any property damage or injury outcomes from SGO-reported crashes, this paper also compares police-reported and any-injury-reported subsets of the SGO data to human benchmarks.
Many previous studies compared early ADS testing data from California, where a human behind the wheel supervised the ADS, sometimes referred to as Testing Operations (TO) (Banerjee et al., 2018; Blanco et al., 2016; Dixit et al., 2016; Favarò et al., 2017; Goodall, 2021; Schoettle and Sivak, 2015; Teoh and Kidd, 2017). All of these early studies use data prior to 2020 when RO operations started and thus the results from the current study may be more relevant in understanding the current state of the art of RO deployments.
Several studies have examined either more recent TO data or RO data. The human benchmark derived by Flannagan et al. (2023) from ride-hailing NDS was compared to the first 1 million miles of RO driving by the company Cruise, which also operated an RO ride-hailing in San Francisco (Zhang, 2023). Zhang (2023) found a 65% reduction in the any property damage or injury outcome category, compared to a 70% reduction found in the current study (0.30 rate ratio in San Francisco compared to the Flannagan et al. (2023) benchmark). Chen and Shladover (2024) compared both testing operations and RO crash rates in San Francisco to a human ride-hailing driver crash rate collected as part of data reporting required in California by the California Public Utility Commission (CPUC). The reporting requirement for this CPUC human ride-hailing crash data is not clearly specified, but presumably the requirement is close to the any property damage or injury threshold. Chen and Shladover (2024) found the human ride-hailing crash rate was 15.5 IPMM in 2020, compared to an 14.1 IPMM for just under 1 million miles of Waymo RO vehicles from September 2022 to August 2023, which represents the ADS reducing the crash rate by 9% compared to the benchmark. No statistical tests were performed by Chen and Shladover (2024) due to the low mileage in the ADS data. The human ride-hailing benchmark in San Francisco reported by Chen and Shladover (2024) is higher than the Scanlon et al. (2023) Blincoe-adjusted any property damage or injury benchmark and lower than the Flannagan et al. (2023) ride-hailing NDS benchmark. Due to an unclear lower reporting threshold in the CPUC data used to derive the human ride-hailing benchmark in Chen and Shladover (2024), it is difficult to interpret the discrepancies between the Chen and Shladover (2024) and Flannagan et al. (2023) benchmarks because the unclear reporting mechanisms in the CPUC data and their implications on underreporting in the human benchmark.
Cummings (2024) compared SGO-reported collisions in California from 2021 to 2022 to a police-reported crash rate. The data used in Cummings (2024) mixed RO and TO data from California. Cummings (2024) compared a Waymo ADS any property damage or injury crash (all SGO-reported crashes) to a police-reported crash rate derived from national crash and VMT data. As discussed previously, the SGO reporting threshold includes any property damage with no lower reporting threshold, whereas police-reported data requires higher amount of property damage or an injury. Therefore, comparing the ADS SGO crash rate to a police-reported crash rate introduces selection bias that deflates the human benchmark relative to the ADS data. Furthermore, the effects of underreporting in the human data were not accounted for in Cummings (2024), introducing reporting bias that also deflates the human crash rate relative to the ADS crash rate. Due to these biases, Cummings (2024) incorrectly concluded that the Waymo ADS had a statistically significant 4-times-higher crash rate than the human benchmark. Indeed, when comparing the Waymo RO any property damage or injury rate in San Francisco from Cummings (2024) (16.5 IPMM) to the national Scanlon et al. (2023) observed police-reported benchmark (4.10 IPMM), the ADS mistakenly appears to have a 4.0-times-higher crash rate. The any property damage or injury and police-reported comparisons done in the current paper control for these selection and reporting biases by using benchmarks that attempt to match the reporting threshold and account for underreporting, and additionally use benchmarks restricted to passenger vehicles and from the areas the ADS operates in (the police-reported crash rate in San Francisco of 5.86 IPMM is 40% larger than the national average of 4.10 IPMM).
The benchmarks and ADS data presented in this paper are overall crash involvement rates which do not consider crash contribution of those involved. Di Lillo et al. (2023) compared third party liability property damage and bodily injury insurance claims rates between the Waymo ADS operations and a human benchmark. A liability claim is a request for compensation when someone is responsible for damage to property or injury to another person, typically following a collision. In Di Lillo et al. (2023), only 3rd party liability claims that were resolved or were likely to be resolved with a payment were included as these types of claims are a proxy for responsibility. Di Lillo et al. (2023) found that over 3.8 million miles of RO driving, the Waymo RO service had a statistically significant, lower property damage and bodily injury claims rate compared to a human benchmark based on aggregate human insurance claims data restricted to zip codes where the Waymo service operates. Examining third party liability claims allows an analysis of the degree to which the ADS contributes to crash causation, that is, the ADS safety impact on other road users. The results of the current study, in contrast, examine the overall crashed vehicle rate regardless of the contribution to the crash’s causation. Many of the crashes in the RO dataset examined in this paper featured the Waymo ADS parked or stopped appropriately at a traffic control device where the ADS’s behavior had little or no contribution to the crash’s cause. Examining ADS performance from both a crash causation and overall crash rate is important because it provides assessment of (a) performance in crashes where there was some contribution (3rd party liability claims), (b) performance in crashes where there was no contribution, and (c) the combination (overall crash rates). Di Lillo et al. (2023) reported that Waymo RO operations reduced property damage liability claims by 76% and bodily injury liability claims by 100%. This magnitude of claims reduction is similar to the magnitude in police-reported (49% to 71% reduction) and any-injury-reported (76% and 90% reduction) reductions reported in the current study, which did not consider contribution to causing the crash.
Interpretation of Results
Any Property Damage or Injury
The available benchmark crashed vehicle rates for any property damage or injury crashes vary widely depending on the method of estimation (i.e., underreporting adjustment of crash data vs. observations in NDS) and the driving environment (e.g., SF ride-hailing NDS vs. national average NDS). Both NDS benchmarks used a four-level severity scale from the SHRP-2 NDS (Hankey et al., 2016) and include level 1 through level 3 severity crashes in their benchmark Flannagan et al. (2023); Blanco et al. (2016). In particular, the most frequent level 3 events include events like road departures and curb strikes with an “increased risk element” without vehicle damage, as well as crashes resulting in minor damage. This level 3 definition does not match the SGO-reporting requirement of property damage nor does it likely match the various crash definitions used to construct the underreporting adjustment in Blincoe et al. (2023). This may partially explain why the NDS crash rates are higher than the Blincoe-adjusted rates. Even still, given the available data it is unclear what magnitude of effect on the any property damage or injury benchmark is due to unclear or different reporting thresholds, underreporting differences by location, or whether the different characteristics of the driving populations (all drivers vs ride-hail drivers using rental vehicles). For example, Scanlon et al. (2023) found that state data from California had a much lower proportion of reported property damage only crashes relative to injury crashes compared with Arizona and national data, which strongly suggests the national Blincoe et al. (2023) may not generalize to individual states. Blanco et al. (2016) attempted to estimate the rate of collisions that met a reporting threshold but were not reported to police, which resulted in a 84% property damage only underreporting amount compared to a 60% underreporting found by Blincoe et al. (2023). Although the Blanco et al. (2016) study had fewer participants and other limitations, the discrepancy suggests there is potential variability in the property damage underreporting amount.
A conservative interpretation of the results of the any property damage or injury outcome group is that conclusions cannot be drawn for this outcome group until benchmarks are further developed. The comparisons in the current study do show, however, the possibility of the ADS reducing crash rates. The ADS reduced crash rates in the comparison of all SGO crashes to a Blincoe-adjusted police-reported rate in Phoenix at a statistically significant level, but a higher (but not statistically significant) crash rate for the Blincoe-adjusted rate in San Francisco. The lower property damage only reporting in California discussed above suggests that the true crash rate in San Francisco may be higher than was estimated from the police-reported crash data, but to what degree could not be determined. The comparisons of all SGO collisions to the NDS benchmarks from Flannagan et al. (2023) and Blanco et al. (2016) showed a large, statistically significant reduction in crash rate for the ADS. As discussed above, however, there is likely events that do not result in property damage in the NDS data samples (i.e., the “level 3” events from the SHRP-2 event definitions), and thus the NDS crash rates are likely inflated to some degree relative to the ADS crash rate. The comparison of SGO collisions excluding crashes with delta-V less than 1 mph showed that many of the ADS collisions are of low severity. As a result this comparison showed the ADS had a reduction in crash rate at a statistically significant level in all locations except for San Francisco (which again could be partially explained by the property damage reporting discrepancies in California).
@@ DISCUSSION
To address these discrepancies and difficulties for the any property damage or injury benchmarks, there is a need for future research that can more accurately define an objective lower reporting threshold in human crash data than exists today. Such benchmarks would also require standardized methods for reporting the severity of ADS crashes, like the delta-V threshold used in this study. The availability of sensor data enables accurate reconstructions of ADS crashes to estimate their severity (e.g., delta-V or impact speed). Estimating crash severity is often not possible from most human crash databases without detailed reconstructions. There are challenges in generating benchmarks for a specific driving environment using sampled or non-representative database that do have such crash reconstructions. Another potential source of severity benchmark is to estimate severity from NDS data. Advances in computer vision may enable crash reconstructions from limited vehicle sensor and video data (see for example Campolettano et al. (2023)).
Police-Reported and Any-Injury-Reported
@@
This comparison of police-reported crashed vehicle rates is likely a conservative comparison for several reasons. Private citizens often do not report police-reportable events (M. Davis & Co., 2015). In California, only injury crashes are required to be reported to the state databases, which means some jurisdictions in California may have different PDO police-reported crashes reporting practices. Conversely for ADS crashes, this study conservatively used the “Law Enforcement Investigating?” field of the NHTSA SGO reports because the authors have knowledge that Waymo’s operational policies in which it is likely police took information that could be used to file a report or indicated they would file a report. This field, as Waymo reports to the SGO, is only reported “yes” when law enforcement indicate they will file a report, not only when law enforcement responds to the scene of a crash. Although the police-reported outcome is often associated with a certain level of property damage, factors other than property damage may influence whether an ADS fleet operator report crashes to the police. Future research should investigate police reporting differences between businesses (like an ADS fleet operator) and the general population.
@@
The Any-injury-reported outcome group may also serve as a useful comparison between ADS and a benchmark. It has been shown that there is less underreporting in injury crashes compared to crashes with no injuries Blincoe et al. (2023). Because the magnitudes of underreporting are smaller, the effect of the underreporting estimate on the comparisons may also be lower, as shown by the results of this study. The injury rates, however, can also suffer from similar surveillance biases as the property damage and police report data related to underreporting. The minimum threshold of an injury is often unclear and is dependent on self-reported injury data. Most injuries in the ADS dataset were minor and did not require medical treatment. As more RO miles are driven, and rarer injury crashes that require medical treatment occur, there is an opportunity to use coded injuries scales, like the Abbreviated Injury Severity (AIS) (AAAM, 2016), to more precisely define injury thresholds.
Other Limitations
@@
In addition to the limitations discussed above, this study has several additional limitations. The results in this study examine crashed vehicle rates aggregated by operating location and do not attempt to evaluate crashed vehicle rates within subgroups, for example, by crash partner or crash type. Past studies that reported ADS crash data have found that ADS vehicles may have different proportions of different crash types than the general population (Victor et al., 2023; Teoh and Kidd, 2017). A lower aggregate crash rate does not preclude an ADS from having an elevated crash rate in certain types of crashes. As more RO data is collected and statistical power allows, future research should investigate crash rates of more granular crash types.